Consider a Hall-effect sensor using 50.0-μm-thick silicon, doped to make it an N-type semiconductor with free electron density 2.86×10^(15) electrons / cm^(3)

With a sensor carrying 550-μA

What is the maximum Hall potential at a point where the earths Magnetic field is 28.5 μT
Answer #1

we know,

Hall potential, VH = B*I/(n*e*d)

given
d = 50.0 micro m
n = 2.86*10^15 electrons/cm^3

= 2.86*10^21 electrons/m^3

e = 1.6*10^-19 C
I = 550 micro A = 550*10^-6 A
B = 28.6*10^-6 T

so,

VH = 28.6*10^-6*550*10^-6/(2.86*10^21*1.6*10^-19*50*10^-6)

= 6.88*10^-7 V (or) 0.688 micro V <<<<<<------------------Answer